# Unveiling Impact of Frequency Components on Membership Inference Attacks for Diffusion Models

### v2

@@ 5.3 In-depth Analysis of Attack Performance
We have theoretically proven that removing high-frequency deficiency can amplify the distinction between member and hold-out data. To further validate our conjectures, we visualize the membership scores of the baselines before and after applying the filter. As illustrated in Fig. 2, it is evident that the distribution gap between member data and hold-out data has increased noticeably after applying our method. Blue boxes mark the areas where member and hold-out data interleave, indicating member/hold-out data are indistinguishable by thresholds. After applying the filter, we observe a significant reduction in sample interleavements. This compelling evidence validates our conjectures and demonstrates the filter’s effectiveness. More visualizations will be presented in Appendix D.6.
@@ 5.4 Ablation Study
To investigate the impact of the filter under different hyperparameter settings, we adjusted the high-frequency threshold $r_{t}$ and the filtering parameter $s$. Experiments were conducted with various values for $r_{t}$ and $s$ on the MS-COCO dataset using Naive attack, and the results are presented in Tab. 4. From the experimental analysis, we recommend a value range for $r_{t}$ of $[3,10]$ and for $s$ of $[0.0,0.3]$. Within this range, the filter achieves optimal performance, significantly enhancing the baseline performance while exhibiting low sensitivity to changes in hyperparameters, further demonstrating the robustness of the filter. When $r_{t}=1$, the high-frequency threshold is set too low, leading to most frequency components of the image being filtered. This contradicts our intention to only suppress high-frequency deficiency, resulting in a significant decline in attack performance. When $s=0.4$ and $s=0.5$, the filtering effect on the high-frequency components is weakened. Although the baseline performance improves significantly, it does not reach the optimal state. More ablation experiments will be presented in Appendix D.8.
Table 4: Naive’attack performance at different $s$ and $r_{t}$ in MS-COCO dataset. Our method has lower sensitivity to hyperparameters and performs excellently under most parameter settings.
$s/r_{t}$
$s=0.0$ $s=0.1$ $s=0.2$ $s=0.3$ $s=0.4$ $s=0.5$
ASR AUC ASR AUC ASR AUC ASR AUC ASR AUC ASR AUC
$r_{t}=3$ 93.70 97.63 93.59 97.59 92.69 96.47 89.20 94.22 84.79 91.93 83.39 90.04
$r_{t}=5$ 94.10 97.89 94.59 97.85 93.60 98.32 90.60 95.23 87.59 93.61 84.89 91.50
@@
Weaker Overfitting. When fine-tuning stable diffusion, the number of iterations is typically determined by the user’s specific needs. In some scenarios, extensive iterations are not required to fine-tune the model. To simulate a model with a lower degree of overfitting, we halved the number of iterations across all datasets, as referenced in Zhai et al. (2024). As shown in Tab. 5, the filter demonstrates nice effectiveness. Taking the Flickr dataset as an example, after applying the filter, the ASR improved by an average of 4.34%, with a maximum improvement of 9.25% in Naive. The AUC increased by an average of 7.04%, with the Naive demonstrating the most significant improvement of 14.53%. Compared to the default settings, the filter’s effect is diminished, which is directly linked to the performance decline of the baseline under the weaker overfitting. Therefore, the results can still prove the effectiveness of the high-frequency filter.
@@
Method
Pokémon MS-COCO Flickr
@@ 5.6 Impact of Defense
$SS_{e_{i}}$ DataAug Naive Naive+F Gain SecMI SecMI+F Gain PIA PIA+F Gain
$\times$ $\times$ 87.95 96.81 +8.86 86.67 91.42 +4.75 82.64 92.54 +9.90
$\times$ $\checkmark$ 86.87 94.14 +7.27 83.16 89.73 +6.57 76.76 85.44 +8.86
$\checkmark$ $\times$ 70.23 76.91 +6.68 81.00 86.82 +5.82 68.81 76.32 +7.51
$\checkmark$ $\checkmark$ 69.09 75.23 +6.14 79.32 84.42 +5.10 61.09 67.95 +6.86
@@ 6 Conclusion
Limitations. Our method exhibits limited efficacy in the pre-training setting, likely due to the overall substandard performance of existing attacks in this setting. This issue may arise from the fact that the pre-trained model does not align with the current assumptions related to overfitting. Future investigations need to thoroughly explore MIAs in the pre-training setting.
@@ A.1 Denoising Diffusion Probabilistic Model (DDPM)
Wen et al. (2024) identified that when a model exactly memorizes training data, the noise prediction network exhibits a pronounced discrepancy between its conditional and unconditional predictions. Given a training data $x$, and the caption embedding $e$ consisting of $N$ tokens, they formulate the minimization objective at step $t$ as:
@@ Appendix B General Paradigm
Denote by $\bm{\Phi}_{\theta}(x_{s},t)$ the deterministic reverse, i.e., from $x_{s}$ to $x_{t}$ ($s<t$):
$x_{t}=\bm{\Phi}_{\theta}(x_{s},t)=\phi_{\theta}(\cdots\phi_{\theta}(\phi_{\theta}(x_{s},s),s+1),t-1),$ (B.14)
and $\bm{\Psi}_{\theta}(x_{t},s)$ the deterministic denoise process, i.e., from $x_{t}$ to $x_{s}$:
$x_{s}=\bm{\Psi}_{\theta}(x_{t},s)=\psi_{\theta}(\cdots\psi_{\theta}(\psi_{\theta}(x_{t},t),t-1),s+1).$ (B.15)
@@
where $\tilde{x}_{t}=\bm{\Phi}_{\theta}(x_{0},t)$. A series of transformations are designed to fetch $x_{t+1}^{target}$ and $x_{t}^{target}$ and predict $x_{t}$ by $x_{t+1}^{target}$. Therefore, Eq. B.16 can be expressed as:
@@
Appendix C Proof of Proposition 1
Proposition 1.
Denote the original standard deviations of membership scores in member and hold-out data as $\sigma_{M}$ and $\sigma_{H}$, and the standard deviations after removing the high-frequency components are $\sigma_{M}^{\prime}$ and $\sigma_{H}^{\prime}$. The standard deviation of membership scores in the high-frequency components is $h_{M}(h_{H})$, and the low-frequency components is $l_{M}(l_{H})$ in member(hold-out) data. Let $l_{H}-l_{M}=\Delta$, $h_{M}=k\cdot h_{H}$ with $k>0$. If $k^{2}>1+\frac{2\Delta}{h_{H}^{2}}(l_{M}+2\Delta-\sqrt{(l_{M}+2\Delta)^{2}+h_{H}^{2}})$, we have:
@@ Appendix C Proof of Proposition 1
The standard deviation before and after filtering high-frequency information satisfies $\sigma_{H}>\sigma_{M}$ and $\sigma_{H}^{\prime}>\sigma_{M}^{\prime}$. $\frac{\sigma_{H}^{\prime}}{\sigma_{M}^{\prime}}>\frac{\sigma_{H}}{\sigma_{M}}$ is equivalent to:
@@
Let $\sigma_{H}=\sigma_{H}^{\prime}+\Delta_{H}$ and $\sigma_{M}=\sigma_{M}^{\prime}+\Delta_{M}$. Substituting these into Eq. C.2, we obtain:
@@
Since $\sigma_{H}^{\prime}>\sigma_{M}^{\prime}>0$, Eq. C.3 holds constantly if and only if:
@@
Let the errors of the high- and low-frequency are independent, denoted as $cov(l_{M},h_{M})\approx cov(l_{H},h_{H})\approx 0$. From the principle of normal distribution superposition, this gives:
$\Delta_{M}=\sqrt{l_{M}^{2}+h_{M}^{2}+cov(l_{M},h_{M})}-\sqrt{l_{M}^{2}},\Delta_{H}=\sqrt{l_{H}^{2}+h_{H}^{2}+cov(l_{H},h_{H})}-\sqrt{l_{H}^{2}}.$ (C.5)
Eq. C.4 can be expressed as:
$\sqrt{l_{M}^{2}+h_{M}^{2}+cov(l_{M},h_{M})}-\sqrt{l_{H}^{2}+h_{H}^{2}+cov(l_{H},h_{H})}>l_{M}-l_{H}.$ (C.6)
Since $l_{H}-l_{M}=\Delta=\sigma_{H}^{\prime}-\sigma_{M}^{\prime}>0$ and $cov(l_{M},h_{M})\approx cov(l_{H},h_{H})\approx 0$, we obtain:
$\sqrt{l_{M}^{2}+h_{M}^{2}}>\sqrt{(l_{M}+\Delta)^{2}+h_{H}^{2}}-\Delta.$ (C.7)
Square both sides of Eq. C.7 to obtain the following:
Substitute $h_{M}=k\cdot h_{H}$ into Eq. C.8, then:
Add $(l_{M}+\Delta)^{2}+h_{H}^{2}$ on both sides of Eq. C.9:
Let $t=(l_{M}+\Delta)^{2}+k^{2}h_{H}^{2}$, substitute $t$ into Eq. C.10:
@@
$(\sqrt{t}+\Delta)^{2}>(l_{M}+2\Delta)^{2}+h_{H}^{2}.$ (C.12)
Therefore, we have:
$\sqrt{t}>\sqrt{(l_{M}+2\Delta)^{2}+h_{H}^{2}}-\Delta.$ (C.13)
Substitute back $t=(l_{M}+\Delta)^{2}+k^{2}h_{H}^{2}$ to Eq. C.13, we obtain:
We ultimately obtain:
Therefore, Proposition 1 is proved. ∎
@@ D.2 Discussion of the theoretical analysis
Constraint Conditions. We conducted experimental validation of the constraint in Proposition 1. As shown in Tab. 7, the parameter $k$ satisfies the constraint across different datasets and models, and in many cases, it even holds that $k>1$. This further demonstrates the practical feasibility and effectiveness of our theoretical results.
@@
Method
Tiny-IN STLU-10 MS-COCO Flickr
@@
Method
MS-COCO Flickr
@@
Method
MS-COCO Flickr
@@ D.5 The Generality of High-Frequency Filtering
The baseline methods (Naive, SecMI, and PIA) rely on the reconstruction errors of images when calculating membership scores, and our method significantly improves their performance. In fact, any attack that relies on the model’s reconstruction capability can consider employing our method. We also add experimental results for CLID (Zhai et al., 2024), which is based on likelihood estimation. Testing under the assumption of weak overfitting as outlined in Sec. 5.5. As shown in Tab. 12, it demonstrates that our method remains effective and shows significant performance improvements across different datasets. Overall, any methodology leveraging the model’s capabilities for image processing can benefit from the frequency analysis and the techniques for mitigating high-frequency deficiency presented in this paper, as this flaw is inherently present in diffusion models.
Table 12: Attack performance of CLID (Zhai et al., 2024) under weak overfitting conditions.
Method
Pokémon MS-COCO Flickr
@@ D.7 Missing Text
Method
MS-COCO Flickr
@@
D.8 Ablation Study on DDIM
As shown in Tab. 15, ablation experiments were performed on DDIM using the Tiny-IN dataset, and the experimental phenomenon is similar to the results in fine-tuned stable diffusion. we recommend a value range for $r_{t}$ of $[3,10]$ and for $s$ of $[0.0,0.3]$. Within this range, the filter achieves optimal performance, significantly enhancing the baseline performance while exhibiting low sensitivity to changes in hyperparameters.
Table 15: Naive’attack performance at different $s$ and $r_{t}$ in Tiny-IN dataset.
$s/r_{t}$
$s=0.0$ $s=0.1$ $s=0.2$ $s=0.3$ $s=0.4$ $s=0.5$
@@ D.8 Ablation Study on DDIM
In addition, we visualize ASR and AUC with different parameter settings. As illustrated in Fig. 7 our method demonstrates extreme robustness and is highly insensitive to hyperparameter variations.
@@
D.9 ROC curves Visualization
As illustrated in Fig. 8, We visualized the ROC curves for different attacks on MS-COCO and Tiny-IN datasets. The blue curves and green curves represent the baselines before and after applying the high-frequency filter module. We can clearly see the powerful effect of the filter through the ROC curve. At the same time, we visualize the ROC curves for different Naive parameter settings on MS-COCO and Tiny-IN datasets. The results are shown in Fig. 9, which show that our method provides a significant improvement over the baselines at different parameter settings.
(a) Naive-MS-COCO
(b) PIA-MS-COCO
(c) SecMI-MS-COCO
(d) Naive-CIFAR-100
(e) PIA-CIFAR-100
(f) SecMI-CIFAR-100
Figure 8: ROC curves before and after applying the high-frequency filter for the baselines.
(a) Naive-MS-COCO
(b) Naive-CIFAR-100
Figure 9: ROC curves of Naive with different parameter settings.
D.10 Compute Overhead

### v3

@@ 5.3 In-depth Analysis of Attack Performance
To further validate that removing high-frequency artifacts can amplify the distinction between member and hold-out data, we visualize the membership scores of the baselines before and after applying the filter. As illustrated in Fig. 2, it is evident that the distribution gap between member data and hold-out data has increased noticeably after applying our method. Blue boxes mark the areas where member and hold-out data interleave, indicating member/hold-out data are indistinguishable by thresholds. After applying the filter, we observe a significant reduction in sample interleavements. This compelling evidence validates our conjectures and demonstrates the filter’s effectiveness. More visualizations will be presented in Appendix D.6.
Table 4: Naive’attack performance at different $s$ and $r_{t}$ in MS-COCO dataset. Our method has lower sensitivity to hyperparameters and performs excellently under most parameter settings. Bold indicates performance using our adopted hyperparameters.
$s/r_{t}$ $s=0.0$ $s=0.1$ $s=0.2$ $s=0.3$ $s=0.4$ $s=0.5$
ASR AUC ASR AUC ASR AUC ASR AUC ASR AUC ASR AUC
$r_{t}=3$ 93.60 97.63 93.59 97.59 92.69 96.47 89.20 94.22 84.79 91.93 83.39 90.04
$r_{t}=5$ 93.50 97.89 93.59 97.85 93.60 98.32 90.60 95.23 87.59 93.61 84.89 91.50
@@ 5.4 Ablation Study
To investigate the impact of the filter under different hyperparameter settings, we adjusted the high-frequency threshold $r_{t}$ and the filtering parameter $s$. Experiments were conducted with various values for $r_{t}$ and $s$ on the MS-COCO dataset using Naive attack, and the results are presented in Tab. 4. Our method achieves the best performance at $s=5$ and $r_{t}=0.2$. From the experimental analysis, we recommend a value range for $r_{t}$ of $[3,10]$ and for $s$ of $[0.0,0.3]$. Within this range, the filter achieves optimal performance, significantly enhancing the baseline performance while exhibiting low sensitivity to changes in hyperparameters, further demonstrating the robustness of the filter. When $r_{t}=1$, the high-frequency threshold is set too low, leading to most high- and low-frequency components of the image being filtered. This contradicts our intention to only suppress high-frequency deficiency, resulting in a significant decline in attack performance. When $s=0.4$ and $s=0.5$, the filtering effect on the high-frequency components is weakened. Although the baseline performance improves significantly, it does not reach the optimal state. More ablation experiments and analyses are presented in Appendix D.8.
@@ 5.5 More Stringent Attack Conditions
Method Pokémon MS-COCO Flickr
@@
$SS_{e_{i}}$ Aug Naive Naive+F SecMI SecMI+F PIA PIA+F
$\times$ $\times$ 87.95 +8.86 86.67 +4.75 82.64 +9.90
$\times$ $\checkmark$ 86.87 +7.27 83.16 +6.57 76.76 +8.86
$\checkmark$ $\times$ 70.23 +6.68 81.00 +5.82 68.81 +7.51
$\checkmark$ $\checkmark$ 69.09 +6.14 79.32 +5.10 61.09 +6.86
Weaker Overfitting. When fine-tuning stable diffusion, the number of iterations is typically determined by the user’s specific needs. In some scenarios, extensive iterations are not required to fine-tune the model. To simulate a model with a lower degree of overfitting, we halved the number of iterations across all datasets, as referenced in (Zhai et al., 2024). As shown in Tab. 5, the filter demonstrates nice effectiveness. Taking the Flickr dataset as an example, after applying the filter, the ASR improved by an average of 4.34%, with a maximum improvement of 9.25% in Naive. The AUC increased by an average of 7.04%, with the Naive demonstrating the most significant improvement of 14.53%. Compared to the default settings, the filter’s effect is diminished, which is directly linked to the performance decline of the baseline under the weaker overfitting. Therefore, the results prove the effectiveness of the high-frequency filter.
@@ 5.6 Impact of Defense
Adaptive Defense. We propose an adaptive defense strategy that applies random amplitude suppression, reducing the low-frequency components of images to 70~90% of their original values, and adds noise with a variance of $0.05$ to the low-frequency domain during the fine-tuning phase. The core mechanism of this strategy lies in moderately suppressing the representation weights of low-frequency components, thereby reducing the model’s overfitting to low-frequency information in images. Experimental results indicate that adaptive defense methods can mitigate the effectiveness of our approach to some extent. More detailed results are provided in the Appendix D.10.
@@ A.1 Denoising Diffusion Probabilistic Model (DDPM)
(Wen et al., 2024) identified that when a model exactly memorizes training data, the noise prediction network exhibits a pronounced discrepancy between its conditional and unconditional predictions. Given a training data $x$, and the caption embedding $e$ consisting of $N$ tokens, they formulate the minimization objective at step $t$ as:
@@ Appendix B General Paradigm
Denote by $\boldsymbol{\Phi}_{\theta}(x_{s},t)$ the deterministic reverse, i.e., from $x_{s}$ to $x_{t}$ ($s<t$):
$x_{t}=\boldsymbol{\Phi}_{\theta}(x_{s},t)=\phi_{\theta}(\cdots\phi_{\theta}(\phi_{\theta}(x_{s},s),s+1),t-1),$ (B.14)
and $\boldsymbol{\Psi}_{\theta}(x_{t},s)$ the deterministic denoise process, i.e., from $x_{t}$ to $x_{s}$:
$x_{s}=\boldsymbol{\Psi}_{\theta}(x_{t},s)=\psi_{\theta}(\cdots\psi_{\theta}(\psi_{\theta}(x_{t},t),t-1),s+1).$ (B.15)
@@
where $\tilde{x}_{t}=\boldsymbol{\Phi}_{\theta}(x_{0},t)$. A series of transformations are designed to fetch $x_{t+1}^{target}$ and $x_{t}^{target}$ and predict $x_{t}$ by $x_{t+1}^{target}$. Therefore, Eq. B.16 can be expressed as:
@@
Appendix C Proof of Proposition 4.2
Proposition 4.2.
Assuming the attack has a membership advantage $Adv^{M}(\mathcal{A})$. Denote the original standard deviations of its membership scores in member and hold-out data as $\sigma_{M}$ and $\sigma_{H}$. And the standard deviations after removing the high-frequency components are $\sigma_{M}^{\prime}$ and $\sigma_{H}^{\prime}$. The standard deviation of membership scores in the high-frequency components is $h_{M}(h_{H})$, and the low-frequency components is $l_{M}(l_{H})$ in member(hold-out) data. Let $l_{H}-l_{M}=\Delta$, $h_{M}=k\cdot h_{H}$ with $k>0$. If $k^{2}>1+\frac{2\Delta}{h_{H}^{2}}(l_{M}+2\Delta-\sqrt{(l_{M}+2\Delta)^{2}+h_{H}^{2}})$, we have:
@@
The standard deviation before and after filtering high-frequency information satisfies $\sigma_{H}>\sigma_{M}$ and $\sigma_{H}^{\prime}>\sigma_{M}^{\prime}$. $\sigma_{H}^{\prime}/\sigma_{M}^{\prime}>\sigma_{H}/\sigma_{M}$ is equivalent to:
@@
Let $\sigma_{H}=\sigma_{H}^{\prime}+\Delta_{H}$ and $\sigma_{M}=\sigma_{M}^{\prime}+\Delta_{M}$. Eq. C.2 is equivalent to:
@@
If we have:
Eq. C.4 can be expressed as:
Let $t=(l_{M}+\Delta)^{2}+k^{2}h_{H}^{2}$, we have:
$\sqrt{t}>\sqrt{(l_{M}+2\Delta)^{2}+h_{H}^{2}}-\Delta.$ (C.6)
Therefore, we have:
$(\sqrt{t}+\Delta)^{2}>(l_{M}+2\Delta)^{2}+h_{H}^{2},$ (C.7)
which is equivalent to:
Substitute back $t=(l_{M}+\Delta)^{2}+k^{2}h_{H}^{2}$ to Eq. C.8, we obtain:
which is equivalent to:
Substitute $k=h_{M}/h_{H}$ into Eq. C.10:
Then, we obtain:
$\sqrt{l_{M}^{2}+h_{M}^{2}}>\sqrt{(l_{M}+\Delta)^{2}+h_{H}^{2}}-\Delta.$ (C.12)
Since $l_{H}-l_{M}=\Delta=\sigma_{H}^{\prime}-\sigma_{M}^{\prime}>0$, we obtain:
$\sqrt{l_{M}^{2}+h_{M}^{2}}-\sqrt{l_{H}^{2}+h_{H}^{2}}>l_{M}-l_{H}.$ (C.13)
Since $\sigma_{H}=\sigma_{H}^{\prime}+\Delta_{H}$ and $\sigma_{M}=\sigma_{M}^{\prime}+\Delta_{M}$, the errors of the high- and low-frequency are independent. From the principle of normal distribution superposition, this gives:
$\Delta_{M}=\sqrt{l_{M}^{2}+h_{M}^{2}}-\sqrt{l_{M}^{2}},\Delta_{H}=\sqrt{l_{H}^{2}+h_{H}^{2}}-\sqrt{l_{H}^{2}}.$ (C.14)
Therefore, we have:
@@
So, we have:
$\sigma_{H}^{\prime}\sigma_{M}^{\prime}+\sigma_{H}^{\prime}\Delta_{M}>\sigma_{H}^{\prime}\sigma_{M}^{\prime}+\Delta_{H}\sigma_{M}^{\prime},$ (C.16)
@@
$\sigma_{H}^{\prime}/\sigma_{M}^{\prime}>\sigma_{H}/\sigma_{M}.$ (C.17)
Therefore, Proposition 4.2 is proved. ∎
@@ D.2 Discussion of the theoretical analysis
Constraint Conditions. We conducted experimental validation of the constraint in Proposition 4.2. As shown in Tab. 7, the parameter $k$ satisfies the constraint across different datasets and models, and in many cases, it even holds that $k>1$. This further demonstrates the practical feasibility and effectiveness of our theoretical results.
@@
Method Tiny-IN STLU-10 MS-COCO Flickr
@@
Method MS-COCO Flickr
@@
Method MS-COCO Flickr
@@ D.5 The Generality of High-Frequency Filtering
The baseline methods (Naive, SecMI, and PIA) rely on the reconstruction errors of images when calculating membership scores, and our method significantly improves their performance. In fact, any attack that relies on the model’s reconstruction capability can consider employing our method. We also add experimental results for CLID (Zhai et al., 2024), which is based on likelihood estimation. We included experiments on CLID under the typical strong overfitting setting in Tab. 12, where it achieves exceptionally high performance. In this setting, our method provides some improvement. Testing under the assumption of weak overfitting as outlined in Sec. 5.5. As shown in Tab. 13, it demonstrates that our method remains effective and shows significant performance improvements across different datasets. Overall, any methodology leveraging the model’s capabilities for image processing can benefit from the frequency analysis and the techniques for mitigating high-frequency deficiency presented in this paper, as this flaw is inherently present in diffusion models.
Table 12: Attack performance of CLID under typical conditions.
Method Pokémon MS-COCO Flickr
ASR AUC TPR@1%FPR ASR AUC TPR@1%FPR ASR AUC TPR@1%FPR
CLID 97.75 99.39 77.00 98.50 99.43 90.00 93.25 96.97 84.00
CLID+F 98.25 99.57 77.00 98.62 99.51 92.00 95.50 98.38 90.00
Table 13: Attack performance of CLID under weak overfitting conditions.
Method Pokémon MS-COCO Flickr
@@ D.7 Missing Text
Method MS-COCO Flickr
@@
D.8 More Ablation Study
Ablation Study on DDIM. As shown in Tab. 16, ablation experiments were performed on DDIM using the Tiny-IN dataset, and the experimental phenomenon is similar to the results in fine-tuned stable diffusion. we recommend a value range for $r_{t}$ of $[3,10]$ and for $s$ of $[0.0,0.3]$. Within this range, the filter achieves optimal performance, significantly enhancing the baseline performance while exhibiting low sensitivity to changes in hyperparameters.
Hyperparameter Analysis. Notably, our method consistently achieves remarkable performance at $s=5,r_{t}=0.2$ across diverse settings, ranging from low-resolution images in DDIM to high-resolution ones in Stable Diffusion. Furthermore, the effectiveness of the proposed method remains robust within our recommended hyperparameter range. This demonstrates that our configuration is highly generalizable and applicable to various datasets. Collectively, these findings reinforce our central thesis that high-frequency deficiency inherently constrain attack performance.
Visualization. In addition, we visualize ASR and AUC with different parameter settings. As illustrated in Fig. 7 our method demonstrates extreme robustness and is highly insensitive to hyperparameter variations.
Table 16: Naive’attack performance at different $s$ and $r_{t}$ in Tiny-IN dataset. Bold indicates performance using our adopted hyperparameters.
$s/r_{t}$ $s=0.0$ $s=0.1$ $s=0.2$ $s=0.3$ $s=0.4$ $s=0.5$
@@ D.8 More Ablation Study
D.9 Compute Overhead